OpenStudy (mm3591):

A(r) is a function that gives the area of a circle with radius r. It can be written in equation form as A(r) = 3.14r2. What is the value of A(5)?

OpenStudy (anonymous):

78.5

OpenStudy (anonymous):

and i thing it is r^2 not r2

OpenStudy (anonymous):

A(5) = 3.14(5)2 I presume would be 3.142(5) instead of 3.14(5)2

OpenStudy (anonymous):

In this case for area of circle the function is Pi (r)^2. Therefore your 3.142 * (5)^2 = Your answer

OpenStudy (anonymous):

put r= 5 in A(r) i.e. A(5)= 3.14(5)^2 =3.14x25 =78.5 unit square